a diagonal of a square lies on the line 3x-5y=14. one vertex is at (0,4). Find an equation of the line that contains the other diagonal.
Answer by Alan3354(8126)

a diagonal of a square lies on the line 3x-5y=14. one vertex is at (0,4). Find an equation of the line that contains the other diagonal.
--------------------------
The diagonals of a square are perpendicular, so the 2nd one will have a slope that's the negative inverse of 3x-5y=14.
The slope, m1, of 3x-5y=14 is 3/5, so m2 will be -5/3.
There are an infinite number to choose from, notice it says "an equation."
If we have the 2nd diagonal pass thru the point (4,0), then find the equation:
y-y1 = m2*(x-x1) where (x1,y1) is the point (4,0)
y = (-5/3)*(x-4)
y = (-5/3)x + 20/3